OpenStudy (anonymous):

PLEASE HELP NOT QUITE SURE ON THIS QUESTION Given the equation −4Square root of x minus 3 = 12, solve for x and identify if it is an extraneous solution.

OpenStudy (anonymous):

this is the equation: \[\sqrt[-4]{x-3} = 12\]

Nnesha (nnesha):

\[\huge\rm -4\sqrt{x-3}=12\] first move the -4 to the right side

Nnesha (nnesha):

it's not 4th root right ? is it -4 (sqrt{x-3} ??

OpenStudy (anonymous):

yes you have it right

OpenStudy (anonymous):

@Nnesha @imqwerty

Nnesha (nnesha):

okay then first move the -4 to the right side how would do that ? divide , multiply add , or subtract ?

OpenStudy (anonymous):

you would add

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

now it's -4 times sqrt{x-3} what is opposite of multiplication ?

OpenStudy (anonymous):

division

Nnesha (nnesha):

yes right so divide both side by -4

OpenStudy (anonymous):

x-3=-2?

Nnesha (nnesha):

12/-4 = ??

OpenStudy (anonymous):

-3

Nnesha (nnesha):

\[\sqrt{x-3}=-3\] now take square both sides to cancel out the square root

OpenStudy (anonymous):

so it would be 0

Nnesha (nnesha):

how ??

OpenStudy (anonymous):

wait no -9

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

x-3=-9

Nnesha (nnesha):

\[\huge\rm (\sqrt{x-3})^2=(-3)^2\] when you take even power of NEGATIVE you will ALAWAYS get positive answer

OpenStudy (anonymous):

oh okay

Nnesha (nnesha):

sp (-3)^2 =? it's same as -3 times -3

OpenStudy (anonymous):

positive 9

Nnesha (nnesha):

x-3= 9 solve for x

OpenStudy (anonymous):

and then you would add 3 to get positive 12

OpenStudy (anonymous):

and then would it be extraneous or not extraneous?

Nnesha (nnesha):

okay to check if it's extraneous or not substitute x for 12 in the original equation

Nnesha (nnesha):

if both sides are equal then 12 isn't extraneous

OpenStudy (anonymous):

both sides came out equal so its not extraneous

Nnesha (nnesha):

how did you get equal sides ?

OpenStudy (anonymous):

honestly i don't know

Nnesha (nnesha):

\[-4\sqrt{12-3}=12\] now solve

Nnesha (nnesha):

show me how did you do it ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im not sure how i did it but could you show me how to do it the right way please?

OpenStudy (anonymous):

i just guessed because i wasn't sure how to do it, the 4 on the outside of the sqr root confuses me

Nnesha (nnesha):

okay so we replaced x with 12 right now solve \[-4\sqrt{12-3}=12\] 12-3 and then take square root

Nnesha (nnesha):

seee i knew it remember it's not 4 it's -4

Nnesha (nnesha):

hint: \[-1\cancel{=}1\]

OpenStudy (anonymous):

i got \[\sqrt[16]{9} =12\]

OpenStudy (anonymous):

that 16 isnt an exponent

Nnesha (nnesha):

hmm how did you get 16? 12-3 = ??

OpenStudy (anonymous):

i did -4^2

Nnesha (nnesha):

don't take square \[-4\sqrt{12-3}\] you should substitute x for 12 into the original equation

Nnesha (nnesha):

12-3= 9 right \[-4\sqrt{9}\] now solve

OpenStudy (anonymous):

equals -12

Nnesha (nnesha):

yes so -12 =12 ?

OpenStudy (anonymous):

sorry as you can tell i struggle with algebra

OpenStudy (anonymous):

no so they are extraneous

OpenStudy (anonymous):

the solution is not extraneous

OpenStudy (anonymous):

solution is extraneous right ? @Nnesha

Nnesha (nnesha):

which one is it extraneous or not ? like i said if you get equal sides THEN 12 is not extraneous

Nnesha (nnesha):

-12=12 both sides are equal ?

OpenStudy (anonymous):

no they are not equal so it is extraneous

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

thanks so much for all your help that took awhile but thank you

Nnesha (nnesha):

my pleasure. :=)